OpenStudy (kj4uts):

The monthly cost C of producing x widgets (in dollars) is modeled by the function shown below. C(x) =2500 + 4.8x + 0.002x^2 In this model 2500 represents fixed monthly costs and 4.8 is the individual cost of each unit being produced. The term 0.002x^2 represents additional costs that are significant only when the production level x is large. Such costs might include additional machinery, time-and-a-half pay, etc. If the maximum monthly cost the company can sustain is $10,500, determine the maximum production level. Round your answer down to the next whole number.

OpenStudy (kj4uts):

@Nnesha

OpenStudy (kj4uts):

Do I just plug in the $10,500 into the formula C(x) =2500 + 4.8(10,500) + 0.002(10,500) ^2?

Nnesha (nnesha):

C= cost x= widgets so you should substitute C for $10,500 solve for x

OpenStudy (kj4uts):

So like this then 10,500 =2500 + 4.8x + 0.002x^2

OpenStudy (kj4uts):

10,500 =2500 + 4.8x + 0.002x^2 I always get stuck here: 8,000=4.8x+0.002x^2

Nnesha (nnesha):

you can use the quadratic formula to solve for x